# 2D-TPE: Two-Dimensional Positional Encoding Enhances Table Understanding for Large Language Models

### v2

@@ 5.2. Results
As shown in Table 3, 2D-TPE is superior to baselines across five datasets, particularly on HiTab, where tables are significantly larger than others. The results reveal the importance of effectively leveraging spatial information. Notably, different datasets may require information in distinct dimensions. For example, Row-wise Traversal significantly outperforms other baselines on HiTab. Conversely, Column-wise Traversal excels in three subsets in TURL. However, these baselines are inherently limited by their 1D perception of the table. In contrast, 2D-TPE allows for token-wise selection of the more valuable spatial dimension.
@@ 5.3. Analysis
To gain deeper insights into the effectiveness and mechanics of our proposed 2D-TPE method, we conduct a comprehensive analysis encompassing several key aspects: investigation of its scalability regarding table sizes (§5.3.1), and the validation of its design choices (§5.3.2), its performance on the larger model and additional baselines  (§5.3.3), the router behavior  (§5.3.4), and efficiency (§5.3.5).
@@
Figure 3. Layer-wise analysis of spatial attention distribution in the fine-tuned model for the Locating-Values task. The graph illustrates the proportion of instances where the model prioritizes column-level information (Orange), row-level information (Grey), or exhibits equal focus on both dimensions (Green).
@@ 5.3.1. Size Scaling
To evaluate the robustness and scalability of 2D-TPE, we conducted a series of experiments focused on size scaling on HiTab. The primary motivation behind this experiment was to assess how well different approaches, including 2D-TPE, can handle tables of increasing complexity and size, which is crucial for real-world applications where table dimensions can vary significantly.
Specifically, we systematically expanded the original table in each example by concatenating additional tables on both the left and right sides from other examples. We considered three settings: the original table (i.e., “0+0”), inserting one table on each side (i.e., “1+1”), and inserting two tables on each side (i.e., “2+2”). The questions and answers remain the same in different settings. More details are shown in Appendix A. This approach allows us to observe the performance trends as the table width increases. As shown in Table 4, the average numbers of columns grow from 8.1 in the original dataset to 21.9 and 35.8 in the 1+1 and 2+2 settings, respectively. Correspondingly, the average table lengths (in tokens) increase from 657 to 1,647 and 2,637, respectively.
The results in Table 4 demonstrate the superior performance and scalability of 2D-TPE compared to baselines. While all baselines show a significant drop in accuracy as table size increases, 2D-TPE maintains relatively stable performance. In contrast, the row-wise traversal method, which performed comparably to 2D-TPE in the “0+0” setting, saw a dramatic decline from 57.65% to 27.52% in the “2+2” setting. The circumstances are similar for other approaches. The robust performance of 2D-TPE across different table sizes underscores its potential as a versatile solution for table understanding tasks in diverse settings.
@@ 5.3.2. Ablation Study
5.3.3. Scaling to Larger Langauge Models
To more convincingly demonstrate the effectiveness of 2D-TPE, we replace the base model with Llama-3-8B (Dubey et al., 2024), which supports a maximum context length of 8,096 tokens. Furthermore, we involve TableLlama (7B) (Zhang et al., 2024b) and GPT-4 as additional baselines.
@@ 5.3.3. Scaling to Larger Langauge Models
Table 6. Results (%) on Llama-3-8B. The italic results are directly taken from Zhang et al. (2024b).
@@
5.3.4. Router Weights
We are curious about how 2D-TPE utilizes the two permutation orders. We calculated the ratio of examples where $\sum_{m,h}r^{h}_{m,1}>\sum_{m,h}r^{h}_{m,2}$ in Eq. 7 for each layer of the fine-tuned model on the test set of the Locating-Values task. Supposing that the subscript $1/2$ corresponds to permutation orders of the row-/column-wise traversal, respectively, we can interpret such examples as instances where the model pays more attention to row-level information than column-level information, and vice versa.
Figure 3 illustrates the ratio of focusing more on rows or columns for each layer of the model. We observe that different layers exhibit distinct tendencies, indicating that 2D-TPE enables the model to dynamically adjust its focus between row-wise and column-wise information processing. This adaptive behavior suggests that the model learns to leverage both dimensions of the table structure effectively, depending on the specific requirements of each layer and the nature of the task at hand. In conclusion, this analysis provides valuable insights into how 2D-TPE facilitates a more comprehensive and adaptable approach to table structure perception.
@@ 5.4. Case Study
Appendix C presents several illustrative cases for the proposed proxy tasks and evaluation benchmarks, providing empirical insights into the efficacy and advantages of 2D-TPE.

### v3

@@ 5.2. Results
As shown in Table 3, 2D-TPE is superior to baselines across five datasets, particularly on HiTab, where tables are significantly larger than others. The results reveal the importance of effectively leveraging spatial information. Notably, different datasets may require information in distinct dimensions. For example, Row-wise Traversal significantly outperforms other baselines on HiTab. Conversely, Column-wise Traversal excels in RelExtra. However, these baselines are inherently limited by their 1D perception of the table. Although M-RoPE achieved moderate performance on EntLink, FeTaQA, and ColType, it exhibited the poorest results on the HiTab dataset. This discrepancy may be because answers can solely be derived from the corresponding tables on HiTab, unlike other datasets that may rely on supplementary information from the questions. Consequently, M-RoPE’s ability to comprehend table structures based on questions may be limited. In contrast, 2D-TPE allows for token-wise selection of the more valuable spatial dimension.
@@ 5.3. Analysis
To gain deeper insights into the effectiveness and mechanics of our proposed 2D-TPE method, we conduct a comprehensive analysis encompassing several key aspects: investigation of its scalability regarding table sizes (§5.3.1), and the validation of its design choices (§5.3.2), its performance when based on larger models (§5.3.3), , and efficiency (§5.3.4).
@@ 5.3.1. Size Scaling
To assess the robustness and scalability of 2D-TPE, we conducted size scaling experiments on HiTab. The goal was to determine how well 2D-TPE and other approaches manage tables of increasing complexity and size, crucial for real-world applications with varying table dimensions.
Specifically, we systematically expanded each original table by adding additional tables to both sides, considering three configurations: the original “0+0”, one table on each side “1+1”, and two tables on each side “2+2”, with unchanged questions and answers. More details are in Appendix A. This method tracks performance as table width grows. As Table 4 shows, average column numbers increased from 8.1 to 21.9 and 35.8, and average table lengths (in tokens) rose from 657 to 1,647 and 2,637, respectively.
Table 4 shows 2D-TPE’s superior performance and scalability over baselines. While all baselines show a significant accuracy drop with larger tables, 2D-TPE remains relatively stable. In contrast, the row-wise traversal method, initially comparable to 2D-TPE, dropped dramatically from 57.65% to 27.52% in the “2+2” setting. Similar trends are observed with other methods. 2D-TPE’s consistent performance across various table sizes highlights its versatility for diverse table understanding tasks.
Furthermore, Figure 3 illustrates the growing advantage of 2D-TPE over three representative baselines as the number of rows or columns increases across three realistic datasets (see Appendix A for full dataset and baseline results). For instance, on FeTaQA, Row-wise Traversal exhibits a mere 0.15% decrease in BLEU-4 score compared to 2D-TPE when the table comprises fewer than 4 columns. However, this gap amplifies to a substantial 2.40% as the number of columns exceeds 8. Similarly, the M-RoPE baseline on the ColType dataset initially lags behind 2D-TPE by 1.84% in the F1 score for tables with fewer than 15 rows, but this deficit exacerbates to 3.13% for tables with more than 30 rows, emphasizing the efficacy of 2D-TPE in dealing with larger tables.
(a) HiTab-Row
(b) HiTab-Column
(c) FeTaQA-Row
(d) FeTaQA-Column
(e) ColType-Row
(f) ColType-Column
Figure 3. Performance advantages ($\Delta$) of 2D-TPE over three representative baselines varying with the number of rows or columns. The thresholds for stratifying tables are determined to ensure a balanced distribution of data volumes.
@@ 5.3.2. Ablation Study
5.3.3. Scaling to Larger Models
To more convincingly demonstrate 2D-TPE’s effectiveness, we replace the base model with Llama-3-8B-Instruct (Dubey et al., 2024), supporting up to 8,096 tokens, and maintain settings from §5.1.3. Furthermore, we involve TableLlama (7B) (Zhang et al., 2024c) and GPT-4 as additional baselines. Due to resource limitations, we do not present the results for EntLink and ColType.
@@ 5.3.3. Scaling to Larger Models
Table 6. Results (%) on Llama-3-8B-Instruct. The italic results are directly taken from Zhang et al. (2024c).
@@ 5.4. Case Study
Appendix C presents several illustrative cases for the proposed proxy tasks and evaluation benchmarks, providing empirical insights into the efficacy, principle, and advantages of 2D-TPE.
@@ Appendix A Details for Size Scaling
Moreover, Table 8 provides a detailed display of 2D-TPE outperforming the complete baselines across five datasets as the number of rows or columns increases. Almost all baselines increasingly lag behind 2D-TPE with the growth in rows or columns, demonstrating the scalability and effectiveness of 2D-TPE, particularly in handling larger datasets.
@@ Appendix B Hyper-parameter Sensitivity
Table 8. Results of 2D-TPE surpassing baselines with increasing rows and columns across various datasets. RT represents Row-wise Traversal, CT stands for Column-wise Traversal, and CA denotes Constrained Attention.
Method HiTab ACC (%) $\uparrow$ EntLink ACC (%) $\uparrow$ FeTaQA BLEU-4 (%) $\uparrow$ RelExtra F1 (%) $\uparrow$ ColType F1 (%) $\uparrow$
Row 0-15 15-30 ¿30 0-16 16-24 ¿24 0-15 15-30 ¿30 0-15 15-25 ¿25 0-15 15-30 ¿30
RT 0.28 2.10 5.39 0.48 0.40 3.05 1.48 0.83 0.02 0.64 0.13 0.93 2.61 1.24 0.96
CT 3.50 6.94 18.86 1.86 1.58 0.34 1.57 1.36 2.02 0.03 0.50 1.81 1.74 0.68 2.12
CA 45.10 44.12 48.82 2.54 4.74 0.00 0.97 0.67 2.56 11.78 12.05 16.98 4.79 3.30 4.84
TABBIE 4.48 4.63 11.78 5.75 8.30 15.76 2.72 2.06 3.56 1.86 1.26 3.27 3.15 1.12 2.88
M-RoPE 56.44 50.84 52.19 2.03 2.37 -1.19 1.55 0.33 0.50 1.66 0.20 1.49 1.84 0.05 3.13
Method HiTab ACC (%) $\uparrow$ EntLink ACC (%) $\uparrow$ FeTaQA BLEU-4 (%) $\uparrow$ RelExtra F1 (%) $\uparrow$ ColType F1 (%) $\uparrow$
Column 0-3 3-7 ¿7 0-4 4-7 ¿7 0-4 4-8 ¿8 0-4 4-7 ¿7 0-4 4-8 ¿8
RT 0.93 2.24 1.65 0.00 2.34 2.05 0.15 1.69 2.40 0.34 0.55 1.15 1.72 2.23 3.66
CT 1.87 7.70 8.56 0.42 0.79 4.62 0.78 1.90 2.31 0.17 0.37 0.20 2.41 1.01 1.69
CA 39.25 44.12 48.05 2.94 1.05 2.56 0.83 0.91 1.50 12.38 12.33 11.79 2.96 5.48 3.55
TABBIE 2.80 5.60 6.91 7.76 11.81 10.77 1.87 3.14 1.85 1.90 1.39 3.59 1.50 3.49 2.20
M-RoPE 52.34 50.42 57.66 1.47 0.39 0.51 0.75 1.26 1.34 0.98 1.39 2.11 1.85 1.55 0.43
@@ C.3. Case Study on Evaluation Benchmarks
Figure 6. Analysis of spatial attention distribution in the fine-tuned model for the Counting-Stars task.
C.4. Router Weights
To investigate how 2D-TPE utilizes the two permutation orders, we fine-tuned MiniCPM-2B-SFT on 10,000 4$\times$4 table data in the Counting-Stars task and randomly selected a sample from 2,000 test set entries. The visualization of the table and the question is shown in Figure 6(a). For the 23rd head, at layer 2, “246” allocates a larger proportion of router weights ( specifically 51.17%) to column-wise traversal, focusing on column-level information, as illustrated in the attention map shown in Figure 6(b). Here, “6” primarily attends to “1$\star$”, “9$\star$”, and “4$\star$”, the stars in the same column as “246”. By layer 27, “246” shifts its focus to row-level information by allocating 52.34% of router weights to row-wise traversal. It distributes most of its attention to “9$\star$” and “1$\star$” in the same row, as depicted in the attention map of Figure 6(c).
This case indicates that 2D-TPE enables the model to dynamically adjust its focus between row-wise and column-wise information processing. This adaptive behavior suggests that the model learns to leverage both dimensions of the table structure effectively, depending on the specific requirements of each layer and the nature of the task at hand. In conclusion, this analysis provides valuable insights into how 2D-TPE facilitates a more comprehensive and adaptable approach to table structure perception.